OpenStudy (anonymous):

The velocity of sound on a particular day outside is 331 meters/second. What is the frequency of a tone if it has a wavelength of 0.6 meters?

OpenStudy (anonymous):

velocity = (frequency) (wavelength)

OpenStudy (anonymous):

331 = f * 0.6 f = 331/0.6